Find two positive integers whose sum is 48, and whose sum also equals twice their difference. What is the smaller and lager integer?
Answer by xcentaur(357)

Let the two numbers be x and y
then x+y=48....[1]
and x+y=2(x-y)...[2]


Then we get,
48=2(x-y)
24=x-y


x+y=48
x-y=24
Adding the two equations we get,
2x=72
x=36


x+y=48
so y=48-x=48-36=12


x=36,y=12


-Prabhat